# High Frequency Quoting Under Liquidity Constraints

### v1

@@ Abstract
Quoting algorithms are fundamental to electronic trading systems, enabling participants to post limit orders in a systematic and adaptive manner. In multi-asset or multi-contract settings, selecting the appropriate reference instrument for pricing quotes is essential to managing execution risk and minimizing trading costs. This work presents a framework for reference selection based on predictive modeling of short-term price stability. We employ multivariate Hawkes processes to model the temporal clustering and cross-excitation of order flow events, capturing the dynamics of activity at the top of the limit order book. To complement this, we introduce a Composite Liquidity Factor (CLF) that provides instantaneous estimates of slippage based on structural features of the book, such as price discontinuities and depth variation across levels. Unlike Hawkes processes, which capture temporal dependencies but not the absolute price structure of the book, the CLF offers a static snapshot of liquidity. A rolling voting mechanism is used to convert these signals into real-time reference decisions. Empirical evaluation on high-frequency market data demonstrates that forecasts derived from the Hawkes process align more closely with market-optimal quoting choices than those based on CLF. These findings highlight the complementary roles of dynamic event modeling and structural liquidity metrics in guiding quoting behavior under execution constraints.
@@ 1 Introduction
1.1 Motivation
High-frequency trading (HFT) has experienced substantial growth over the past decades, significantly influencing modern securities markets. In the U.S., HFT accounted for approximately 20% of equity trading in 2005, which surged to a peak of 60% by 2009 [1]. This rapid expansion is attributed to technological advancements and market participants’ adoption of sophisticated quoting algorithms. These algorithms enable rapid placement and adjustment of buy and sell orders, reacting to real-time market data within microseconds. Such speed and efficiency are essential in today’s liquid markets, where the volume of trading and competing algorithms necessitate swift responses to maintain a competitive edge [2].
Studies have further highlighted the significant role of HFT in price formation and liquidity provision. For instance, Clapham [3] emphasizes that high-frequency traders account for a substantial portion of overall price discovery and liquidity in modern securities markets. Additionally, Brogaard et al. [2] indicate that HFT firms contribute to over 50% of trading volume in U.S. equity markets, underscoring their dominant presence.
More recently, HFT has undergone a significant transformation, shifting from aggressive liquidity-taking strategies to more sophisticated quoting algorithms for execution. During the initial phase of HFT adoption, speed was the primary competitive advantage, with traders employing market order-based strategies to capture fleeting arbitrage opportunities. As low-latency infrastructure became widely available and competition intensified, these aggressive approaches faced diminishing returns due to increased transaction costs, market impact, and adverse selection risks. This evolution in market structure and competitive dynamics has prompted a shift toward execution strategies that prioritize efficiency, resilience, and adaptability. Quoting algorithms offer a sustainable and adaptive approach by intelligently placing and managing limit orders based on evolving market conditions. Rather than crossing the spread with market orders, these strategies shift the burden of research to predictive modeling of near term limit order book (LOB) state rather than the engineering objective of decreased latency. Market participants that use quoting algorithms to execute their trades aim to earn the spread by providing liquidity, thereby reducing execution costs and mitigating exposure to adverse selection. By incorporating predictive models on real-time market data, quoting algorithms allow participants to optimize execution quality, improve liquidity provision, and shift away from strategies that rely solely on latency as a competitive edge. Kearns et al. [4] highlighted that aggressive strategies often suffer from substantial execution costs and adverse price impacts, limiting their profitability. This necessitated a shift towards diversified approaches, particularly automated quoting strategies. By acting as market makers, traders using quoting strategies can maintain consistent income and reduce reliance on speed advantages [5].
At their core, quoting algorithms are automated systems that place and update limit orders on one or both sides of the LOB in response to market dynamics. These algorithms balance multiple objectives—such as reducing inventory risk, capturing the bid-ask spread, and maintaining quote competitiveness—while operating under stringent latency and reliability constraints. In the context of HFT, they are designed to react to microstructure-level events in real time, adjusting quoted prices and order sizes with high precision. Their effectiveness depends not only on speed, but also on the ability to anticipate short-term order flow and market conditions, making predictive modeling an essential component of modern quoting strategies [2].
In efficient markets, arbitrage opportunities are typically short-lived and require precise, rapid execution to be profitably exploited. As markets have become more electronic and competitive, the notion of arbitrage has expanded beyond simple price discrepancies to include statistical and structural inefficiencies that may persist over very short time horizons. Quoting algorithms are increasingly used as tools for arbitrage in this context—not to exploit obvious mispricings, but to identify and respond to subtle imbalances in supply and demand across related instruments or timeframes. This shift has laid the groundwork for more complex arbitrage frameworks involving multiple contracts and coordinated quoting logic.
Multi-contract arbitrage strategies further illustrate the evolution of HFT. These strategies exploit price discrepancies across correlated asset classes, leveraging complex instruments like calendar spreads, three-leg butterflies, and iron condors. For example, calendar spreads profit from time decay or volatility differences between futures or options contracts of the same underlying asset with different expiration dates [6]. Similarly, three-leg butterflies balance risk and reward during minimal price movement [7], while iron condors capture premium decay in range-bound markets [8]. Optimization techniques and dynamic rebalancing are crucial for managing such portfolios [9].
When employing a quoting algorithm for multi-contract arbitrage, a critical decision involves determining which contract or subset of contracts should be the reference price while posting limit orders on others to achieve the required spread. Quoting strategies address the inefficiencies of aggressive liquidity-taking execution. By conditionally placing quotes on a subset of assets while using others as references, these strategies aim to capture trader-defined spreads ($S$) while minimizing execution costs. Unlike reacting to market opportunities after they arise, quoting ensures that limit orders are already in place, increasing the likelihood of execution during favorable price movements. Typically, quoting algorithms place limit orders on a subset of contracts, with the price of the limit orders being a function of $S$ and the reference price. Once the limit order is filled, the quoting algorithm executes an aggressive limit order [^1] The National Stock Exchange (NSE) of India, prohibits market orders from being submitted by algorithmic trading systems, requiring all orders to have specified price limits. Consequently, our framework employs aggressive limit orders to ensure order execution within the permissible structure of the market. on the reference contract. This process assumes that the reference price remains stable during the time it takes for the limit order to be filled and the aggressive order to execute. Ensuring stability in the reference price during this time is crucial, as any significant changes could erode the profitability of the arbitrage trade.
In order to effectively control execution risk, market participants require methods to forecast which contract or subset of contracts is least likely to experience price changes during the critical execution interval. In this context, we explore two methods tailored for a fundamental example of multi-contract arbitrage: the futures calendar spread strategy. Calendar spreads, which involve offsetting positions in futures contracts of the same underlying asset with different expiration dates, are particularly sensitive to price dynamics [6]. Effective forecasting of stable reference contracts enhances the reliability and profitability of quoting strategies in these scenarios.
2 Literature and Existing Frameworks
The existing body of work provides frameworks and analyses to address the practical requirements of high-frequency quoting. [10] introduces a framework for optimal market making in an LOB, balancing profitability and inventory risk. [11] analyzes algorithmic trading strategies, including quoting algorithms, in Algorithmic and High-Frequency Trading. Similarly, [12] examines liquidity provision dynamics and quoting strategies’ role in promoting market efficiency. In multi-asset trading, [13] highlights quoting algorithms’ contributions to market efficiency and price discovery across multiple venues. High-frequency quoting strategies have been widely studied for their roles in market-making, arbitrage, and liquidity provision. These strategies target microstructure inefficiencies and short-term price movements while managing risks like slippage and adverse selection. Studies such as [10] and [14] provide optimal quoting frameworks for market-making, focusing on capturing the spread by quoting bid and ask prices simultaneously. In spread trading, [1] discusses exploiting relationships between correlated instruments, such as futures calendar and butterfly spreads. Effective execution algorithms are crucial for managing multi-leg trades efficiently while maintaining desired spreads. Probabilistic modeling techniques, like Hawkes processes, are powerful tools for analyzing market microstructure. [15] introduces these self-exciting point processes to model event clustering, such as trades and quote updates. Their temporal dependency modeling, as shown by [16], is valuable for forecasting order book dynamics. However, the decay kernel’s choice significantly affects performance, motivating systematic optimization of kernel selection, as addressed in this paper. Real-time indicators based on order book data are integral to high-frequency quoting. [17] and [18] demonstrate how features like order book depth and imbalance inform trading decisions. This paper extends such approaches by introducing the Composite Liquidity Factor (CLF), which aggregates order book features into a single actionable metric. Although existing frameworks provide valuable foundations, they often lack adaptability to varying liquidity conditions. This paper bridges probabilistic forecasting with real-time decision-making by comparing the performance of Hawkes-based modeling and the CLF, offering a unified framework for optimizing high-frequency quoting strategies.
3 Contribution of This Paper
This paper introduces a novel framework for integrating liquidity measures and probabilistic modeling into high-frequency trading strategies. The contributions are summarized as follows:
  • Arrival ratio ($\rho$): A high frequency measure derived from multivariate Hawkes processes that models order arrivals to estimate short-term price stability for reference contract selection.
  • Composite Liquidity Factor (CLF): A high frequency liquidity metric derived from order book data, offering traders an actionable tool for decision-making in dynamic market environments.
  • Benchmark for evaluating quoting strategies: We propose a benchmark against which a quoting strategy can be evaluated.
  • Empirical Validation: Using tick-by-tick data from the Indian National Stock Exchange(NSE), we evaluate the proposed methodology’s impact on slippage reduction and spread capture.
4 Execution Structure and Reference Leg Selection Problem
Execution of calendar spread trades typically involves taking opposing positions in two futures contracts on the same underlying asset but with different expiry dates. While the theoretical spread between these contracts reflects cost-of-carry relationships, the realized execution cost is heavily influenced by the microstructure of the limit order book. In practice, algorithmic traders often employ asymmetric quoting strategies, where a limit order is placed on one leg of the spread (the quoting leg), and upon its execution, an aggressive order is immediately fired on the second leg (the reference leg) to complete the trade. This asymmetric execution introduces a time window—referred to as the critical interval—during which adverse price movements on the reference leg can lead to significant deviation from the intended spread. The central problem we address in this paper is the design of a reference leg selection strategy that minimizes expected slippage under such execution risk. This section formalizes the quoting framework, defines key variables, and motivates the need for a dynamic, data-driven approach to reference leg choice.
4.1 Quoting Mechanics and Execution Risk
@@
A futures calendar spread strategy involves trading two futures contracts on the same underlying asset but with different expiration dates. The objective is to profit from changes in the price difference, or spread, between the near-term and far-term contracts. Table 1 presents the bid and ask prices of NIFTY futures contracts expiring in February 2022 and March 2022. The near-term contract (February) typically trades at a lower price, while the far-term contract (March) reflects higher prices, consistent with cost-of-carry models in futures markets.
@@
Using an aggressive liquidity-taking execution strategy, the trader places aggressive limit orders to immediately execute trades by consuming available liquidity at the best bid and ask prices. [^1] Throughout this paper, references to liquidity-taking strategies should be interpreted as involving aggressive limit orders, not market orders, in compliance with NSE regulations. See Remark 1.
Remark 1.
(On the Use of Aggressive Limit Orders)
At the National Stock Exchange (NSE) of India, algorithmic trading systems are prohibited from submitting market orders. Instead, execution is achieved through aggressive limit orders, which are limit orders posted with prices sufficiently favorable to ensure immediate execution against standing book liquidity. Throughout this paper, references to liquidity-taking strategies should be interpreted accordingly.
Suppose the trader anticipates that the spread between the two contracts will narrow over time due to market conditions, such as a reduced cost of carry or convergence in demand.
The trader initiates the calendar spread by:
@@
We adopt the following notational conventions throughout: $p(t)$ denotes the quoted limit order book price of a futures contract as a function of time $t$, with superscripts $b$ and $a$ indicating the bid and ask sides, respectively. The subscripts $c$ and $n,$ represent the current and the next month. $\tilde{p}(t)$ denotes an executed (traded) price at time $t$, distinguishing it from quoted limit order book prices. The trader defines a target spread $S$ at the time of quoting, representing the desired price differential between legs. The realized (executed) spreads at entry and exit are denoted by $\tilde{S}_{\text{entry}}$ and $\tilde{S}_{\text{exit}}$, respectively, reflecting the effective spread captured through actual trades. The net realized profit $\Pi$ is computed as the difference between the entry and exit spreads. We let $t_{0}$ denote the time of spread entry and $t_{1}$ denote the time of spread exit.
@@
This motivates a critical challenge in quoting strategies: identifying which contract or combination of contracts will exhibit the greatest price stability during the execution window, thereby maximizing the probability of achieving the desired spread.
4.2 Execution using quoting
The same spread trade can be executed through a quoting strategy. In this approach, the trader begins with a predefined target spread $S$. Rather than accepting the current market spread as fixed, the trader aims to minimize the cost of entry and maximize the profitability of the exit by selectively quoting one leg of the spread.
To execute the strategy, the price of one contract is designated as the reference price, and a limit order is placed on the other contract at a price derived as a function of the reference price and the target spread $S$. Upon the limit order being filled, a corresponding aggressive limit order (crossing the book) is triggered on the reference contract to complete the spread execution.
@@ 4.2 Execution using quoting
In quoting strategies, however, execution risk is non-zero. This arises because of the latency between the limit order on the quoting leg being filled and the aggressive limit order on the reference leg being fired and executed. During this time window, the reference price may change adversely.
As a result, the realized spread at time $t$, denoted $\tilde{S}(t)$, may differ from the trader’s intended spread $S$. In certain market conditions such as contango, the farther-month futures price may trade below the nearer-month price, leading to a negative target spread. [^2] See Remark 2 for a discussion on contango markets and reverse spreads. The slippage associated with quoting, denoted $\delta_{S}(t)$, is defined as:
@@
In certain market conditions such as contango, the futures price of a farther-month contract may trade below that of a nearer-month contract. This results in a negative intended spread $S$ for calendar spread strategies. The slippage definition $\delta_{S}(t)=|S-\tilde{S}(t)|$ naturally accommodates both positive and negative target spreads.
4.3 Execution using limit orders
Market makers frequently employ quoting strategies that involve placing limit orders on both the bid and ask sides of the limit order book (LOB), aiming to profit from short-term fluctuations in the bid-ask spread—especially during phases of spread expansion or contraction. However, in the context of futures rollover trades, simultaneously quoting both legs of a multi-contract strategy increases execution risk. The resulting fill asymmetry can lead to a significant deviation between the realized spread $\tilde{S}(t)$ and the trader’s intended spread $S$, reflected in higher slippage $\delta_{S}(t)$.
Another critical consideration when designing quoting strategies is whether the objective is to minimize execution risk alone, or whether it must also meet minimum trading volume constraints over time. While the framework presented in this paper can be extended to symmetric quoting across multiple contracts, our focus is specifically on asymmetric multi-contract quoting strategies, in which limit orders are placed only on a subset of contracts and priced using reference prices derived from a disjoint subset.
Impact of Reference Leg Selection on Execution Slippage
This subsection presents a concrete example of how the choice of reference leg—i.e., which contract’s price is used to trigger the market order—can significantly affect the realized spread and slippage in a calendar spread quoting strategy.
We consider two futures contracts on the same underlying: the current month contract (denoted by subscript $c$) and the next month contract (subscript $n$). The trader wishes to enter a spread trade at time $t_{0}$ with a mandated spread of $S(t_{0})=58.5$ INR per contract.
Time Convention:

### v2

# Event-Time Anchor Selection for Multi-Contract Quoting

@@ Abstract
When quoting across multiple contracts, the sequence of execution can be a key driver of implementation shortfall relative to the target spread [6]. We model the short-horizon execution risk from such quoting as variations in transaction prices between the initiation of the first leg and the completion of the position. Our quoting policy anchors the spread by designating one contract ex ante as a reference contract. Reducing execution risk requires a predictive criterion for selecting that contract whose price is most stable over the execution interval. This paper develops a diagnostic framework for reference-contract selection that evaluates this stability by contrasting order-flow Hawkes forecasts with a Composite Liquidity Factor (CLF) of instantaneous limit order book (LOB) shape. We illustrate the framework on tick-by-tick data for a pair of NIFTY futures contracts. The results suggest that event-history and LOB-state signals offer complementary views of short-horizon execution risk for reference-contract selection.
Keywords: High Frequency Trading, Execution, Slippage, Market Microstructure, Liquidity, Quoting
@@ 1 Introduction
Automated trading in electronic limit order books produces a dense, time-stamped order flow that motivates an event-time rather than calendar-time treatment of short-horizon execution risk [12, 1, 16]. Advances in networking, matching-engine design, and co-located infrastructure have enabled rapid order placement and adjustment at microsecond horizons.
In liquid venues, this latency profile is now routine. With increased adoption of low-latency infrastructure across venues, the marginal value of further latency reductions has declined. Consequently, execution has shifted from predominantly aggressive liquidity-taking policies to adaptive quoting policies [27]. Quoting is effective because it is anticipatory: placements and updates are conditioned on short-horizon forecasts of the microstructure state, rather than solely on the current tick or current limit order book state. In practice, quoting systems place and update limit orders on one or both sides of the LOB using estimated near-term trajectories of order flow, queue depth, and price continuity [2, 20]. Design goals include lowering inventory variance, capturing bid–ask spread, and maintaining price competitiveness under latency and reliability constraints. Quoting performance, therefore, depends on forecasting of short-horizon order flow and the probability of adverse selection, both of which influence spreads and trading costs [22, 21]. Accordingly, in this paper we treat predictive modeling as a central component of quoting and use this perspective to motivate the choice of a reference instrument (basis) in multi-contract settings.
No-arbitrage requirements jointly constrain the prices of related contracts, necessitating coordinated pricing and execution across individual legs [7]. Accordingly, we adopt multi-contract arbitrage (see Section 1.3 for a detailed example) in a diagnostic setting: in liquid markets, arbitrage opportunities are short-lived and spreads are thin. Even small reductions in execution cost translate directly into realized P&L [32, 27]. As illustrative structures, calendar spreads, butterflies, and iron condors are routinely used to express short-horizon views under thin spreads and tight constraints [31, 13, 34]. This sensitivity makes reference-leg selection a measurable problem of near-term basis stability, which we evaluate using only information available at the start of each decision interval.
In multi-contract arbitrage, a quoting policy must select a reference leg and price the target leg’s quotes relative to that reference. Operationally, the strategy posts limit orders on the target leg while monitoring a designated reference price. Once the target order fills, the strategy submits a marketable limit order on the reference leg to complete the position. [^1] On the National Stock Exchange (NSE) of India, market orders are disallowed for algorithmic systems; orders must include explicit price limits, therefore, our implementation uses marketable limit orders on the reference leg. This design places orders in advance rather than reacting after opportunities arise, aiming to capture a trader-defined spread while controlling execution cost. In practice, quote prices are functions of the observed mid-price of the target leg and the contemporaneous reference price. The approach presumes that the reference remains sufficiently stable over the fill horizon. Instability in the basis results in deviation from target spread and reduced profitability [6].In this setting, execution risk is amplified because the reference leg establishes the pricing basis while other legs await fills. Our focus is on the information available at decision time: a temporal component based on multivariate Hawkes processes (order-flow clustering and cross-excitation) and a contemporaneous LOB-state component aggregated as a Composite Liquidity Factor (CLF). In this design, the selection of reference-legs is evaluated from two differing information sets.
Remark.
We study reference selection as an informational problem for execution–cost control, not as a prescriptive quoting rule. Two information sets are considered at decision time: (i) temporal interaction in order–flow events, modeled with a multivariate Hawkes process (clustering and cross–excitation), and (ii) contemporaneous LOB state, summarized as a Composite Liquidity Factor (CLF) derived from depth and price discontinuities. Decisions are evaluated in a rolling, walk–forward, out–of–sample design, with each interval’s choice formed from the information available at its beginning. Evaluation is anchored by a hindsight lower bound on deviation from the target spread, which provides a common yardstick for comparing heterogeneous rules.
1.1 Literature and existing frameworks
A large body of literature models optimal quoting under inventory and adverse selection constraints in dealer and electronic markets [29, 22]. In electronic limit order books, Avellaneda and Stoikov [2] formulate a stochastic control model for dynamic bid/ask placement that trades off expected return against inventory risk, while Cartea et al. [14] survey algorithmic trading frameworks that include inventory-aware quoting. Related microstructure work shows how quoting behavior shapes liquidity and price formation [20]. Collectively, these studies support dynamic quotes that respond to inventory and volatility. However, the canonical formulations are predominantly single-asset and do not address the choice of reference instrument in multi-contract quoting.
From single-asset quoting to multi-contract reference-leg selection
Multi-leg spreads (e.g., futures calendar spreads and option combinations) raise execution challenges that single-asset models do not address. Prior work extends optimal control to multi-asset settings, including stochastic programs for arbitrage portfolios [7] and closed-form policies for multi-asset market making with joint inventory control [5]. These studies show potential efficiency gains from coordinated quoting across related instruments. In practice, legs rarely execute simultaneously: quoting policies place limit orders on the non-reference leg(s) and execute the reference leg last with a marketable limit order once those orders fill [1]. This sequential execution creates a short fill horizon in which adverse price moves against the reference cause the realized spread to deviate from the targeted spread. Multi-asset execution frameworks also address cross-asset coordination [6]. However, while multi-contract quoting has been studied, the specific question of sequencing risk under different choices of the reference leg is, to our knowledge, not established in the literature.
A second line of work studies real-time features derived from LOB state. The configuration of the LOB, including depth at best quotes, LOB imbalance, and the spread, contains predictive information about short-horizon price changes [18, 9]. In a Markovian LOB setting, Cont et al. show that transition probabilities depend on volumes at the best quotes and related state variables [18]. In practice, limit order book features are routinely incorporated as decision inputs in high-frequency execution and market-making systems [14, 27]. Several studies propose composite indicators to summarize liquidity conditions by combining depth, spread, and imbalance. Building on this idea, we construct a Composite Liquidity Factor (CLF) that aggregates instantaneous LOB features into a scalar proxy for near-term price stability derived from depth and price discontinuities, and we use it as a real-time decision input in our framework. To our knowledge, prior studies analyze gap statistics and their relation to jumps and impact, but do not aggregate cross-level gaps with depth into a single factor for reference-leg selection in multi-contract quoting [10, 18, 8, 14].
A third line of work models market activity in event time rather than clock time. Auto-regressive Conditional Duration (ACD) models show that trade durations exhibit serial dependence that can be exploited for short-horizon forecasting [19]. Multivariate self-exciting point processes extend this approach to order-flow interactions: Hawkes models capture clustering and cross-excitation between trades and quote updates and have been applied extensively in high-frequency settings [11, 4, 15, 35]. Empirical studies report that incorporating self-excitation improves short-horizon predictions of price changes and liquidity shortfalls [26]. A practical consideration is the specification of the excitation kernels with both parametric and non-parametric estimators being used in practice [33]. In our framework, the Hawkes arrival ratio provides temporal estimates of near-term order-flow intensity and direction, yielding leg-specific forecasts that inform reference-leg selection. Within Hawkes-based LOB models, (i) queue-reactive specifications endogenize excitation via queue-size–dependent terms [43, 30], while (ii) LOB–dependent specifications let the intensity vary with observable LOB co-variates [36, 41].
In summary, to our knowledge and within the scope of this review, the problem of dynamic reference-leg selection for multi-contract spreads remains underdeveloped. To address this gap, we propose a framework that combines a Hawkes-based estimate of imminent relative price stability with a Composite Liquidity Factor (CLF) summarizing contemporaneous LOB state. We evaluate the resulting choices in a rolling, walk-forward, and out-of-sample design.
1.2 Contribution of This Paper
This paper presents a general framework for incorporating order flow and LOB state in quoting policy. Specifically, we compare an arrival ratio based on Hawkes processes with a composite liquidity factor (CLF) to study reference-leg selection in multi-contract quoting. Under this framework, the contributions of this paper are summarized as follows:
  1. 1.
    Arrival ratio ($\rho$): A leg-specific, Hawkes-based estimate of short-horizon order-flow direction and intensity used for reference-leg selection.
  2. 2.
    Composite Liquidity Factor (CLF): A scalar summary of limit order book state (depth and cross-level price gaps) used as a proxy for short-term price stability in reference-leg selection.
  3. 3.
    Benchmark oracle: A deviation from the target spread metric, compared to an infeasible hindsight lower bound (oracle), enabling like-for-like comparison across heterogeneous rules.
1.3 Execution Structure and Reference Leg Selection Problem
Among the many available multi-contract strategies that employ quoting, this study uses the calendar-spread strategy. The calendar spread offers two advantages: it has the minimal number of legs for a multi-contract design (two), and the only variation between the two contracts is the date of expiration on the same underlying. This concentrates execution effects in the traded spread with fewer confounding factors. Execution of calendar spreads involves taking offsetting positions in two futures contracts with different expiries. We define the reference-leg selection problem as choosing, at each decision time, that contract whose short-horizon dynamics will anchor the spread quote while the other leg awaits execution.
Quoting Mechanics and Execution Risk
@@ 1.3 Execution Structure and Reference Leg Selection Problem
In a futures calendar spread, the traded entity is the price difference, or spread, between the near-term and far-term contracts. Table 1 presents bid and ask prices for NIFTY futures expiring in February 2022 and March 2022. In typical carry regimes, near-term contract prices are below the farther-term contract, consistent with the cost-of-carry term structure.
To illustrate the mechanics of a calendar spread, we consider execution under a liquidity-taking setup. The trader submits marketable limit orders to buy the near-term contract at the best ask and to sell the far-term contract at the best bid, achieving immediate fills. With the sign convention that the calendar spread is quoted as far minus near, this entry corresponds to a long-near/short-far position with a positive spread at execution. [^1] Throughout, references to liquidity-taking strategies indicate marketable limit orders (not market orders). See Remark Remark.
Remark.
(On liquidity-taking via marketable limit orders)
@@
In this paper, liquidity-taking execution is implemented exclusively with marketable limit orders (also referred to as “aggressive” limit orders), i.e., limit orders priced to execute immediately against standing best quotes. We do not use market orders. All references to liquidity-taking strategies should be interpreted accordingly.
We adopt the following notational conventions throughout: $p(t)$ denotes the quoted limit order book price of a futures contract as a function of time $t$, with superscripts $b$ and $a$ indicating bid and ask, respectively. Subscripts $c$ and $n$ refer to the near (current) and far (next) month contracts. $\tilde{p}(t)$ denotes an executed (traded) price at time $t$, distinguishing it from quoted prices. We define the calendar spread as $S(t)=p_{n}(t)-p_{c}(t)$ (next minus current). The realized spreads at entry and exit are denoted $\tilde{S}_{\text{entry}}$ and $\tilde{S}_{\text{exit}}$. The net realized profit is $\Pi=\tilde{S}_{\text{entry}}-\tilde{S}_{\text{exit}}$. We let $t_{0}$ denote the time of spread entry and $t_{1}$ the time of spread exit.
Suppose the trader anticipates that, over the intended holding interval, the calendar spread will narrow (for example, owing to temporary changes in cost of carry or relative demand). Under the liquidity-taking setup described above and the same sign convention, the trader enters a long-near/short-far position and plans to exit once the realized spread has tightened by the targeted amount. The trader initiates the calendar spread by:
  1. 1.
@@
Execution using quoting
The same spread trade can be executed through a quoting strategy. In this approach, the trader begins with a predefined target spread $S$. Rather than accepting the current market spread as fixed, the trader aims to minimize the cost of exit and maximize the profitability of the entry by selectively quoting one leg of the spread.
To execute the strategy, the price of one contract is designated as the reference price, and a limit order is placed on the other contract at a price derived as a function of the reference price and the target spread $S$. Upon the limit order being filled, a corresponding marketable limit order (crossing the book) is triggered on the reference contract to complete the spread execution.
@@
In quoting strategies, however, execution risk is non-zero. This arises because of the latency between the limit order on the quoting leg being filled and the marketable limit order on the reference leg being fired and executed. During this time window, the reference price may change adversely.
As a result, the realized spread at time $t$, denoted $\tilde{S}(t)$, may differ from the trader’s intended spread $S$. In certain market conditions such as contango, the farther-month futures price may trade below the nearer-month price, leading to a negative target spread. [^2] See Remark Remark for a discussion on contango markets and reverse spreads. The slippage associated with quoting, denoted $\delta_{S}(t)$, is defined as:
@@
In certain market conditions, such as contango, the futures price of a farther-month contract may trade below that of a nearer-month contract. This results in a negative intended spread $S$ for calendar spread strategies. The slippage definition $\delta_{S}(t)=|S-\tilde{S}(t)|$ naturally accommodates both positive and negative target spreads.
Execution using limit orders
Market makers frequently employ quoting strategies that involve placing limit orders on both the bid and ask sides of the limit order book, aiming to profit from short-term fluctuations in the bid-ask spread, especially during phases of spread expansion or contraction. However, in the context of futures rollover trades, simultaneously quoting both legs of a multi-contract strategy increases execution risk. The resulting fill asymmetry can lead to a significant deviation between the realized spread $\tilde{S}(t)$ and the trader’s intended spread $S$, reflected in higher slippage $\delta_{S}(t)$.
Another critical consideration when designing quoting policies is whether the objective is to minimize execution risk alone, or whether it must also meet minimum trading volume constraints over time. While the framework presented in this paper can be extended to symmetric quoting across multiple contracts, our focus is specifically on asymmetric multi-contract quoting strategies, in which limit orders are placed only on a subset of contracts and priced using reference prices derived from a disjoint subset.
1.4 Impact of Reference Leg Selection on Execution Slippage
We consider two futures contracts on the same underlying: the current-month contract (denoted by subscript $c$) and the next-month contract (subscript $n$). The trader wishes to enter a spread trade at time $t_{0}$ with a target spread of $S(t_{0})=58.5$ INR per contract. We adopt the following time Convention: